a 13 feet ladder is placed 5 feet away from a wall. the distance from the ground straight upto the top of the wall is 13 feet. how far up the wall will the ladder reach?
Answer by checkley79(3341)

You have a right triangle.
a^2+b^2=c^2
5^2+b^2=13^2
25+b^2=169
b^2=169-25
b^2=144
b=sqrt144
b=12 ft. ans.
Proof:
5^2+12^2=13^2
25+144=169
169=169